OpenStudy (anonymous):

Medal for help-A company produced 4,985 gadets at a cost of $2.07 each. The company sold 4,109 of the gadgets for $14.99 each.Which is the closest estimate of the company's profit or loss ? A. $40,000 loss b. $50,000 loss c $60,000 Profit d. $50,000 profit

OpenStudy (anonymous):

@shiraz14 any clue ?

OpenStudy (anonymous):

Wait no I meant d

OpenStudy (anonymous):

Yes

OpenStudy (shiraz14):

Profit = Total sales - Total expenditure = (4109x14.99)-(4985x2.07) ~ (4100x15)-(5000x2) = 61500-10000 = 51500 (a gain of ~51500) The closest estimate would thus be d.

OpenStudy (anonymous):

Cool Ty I got the same